Four coins are tossed. Let Y be the random variable that represents the number tail that will occur. Find the values of random variable Y.


1
Expert's answer
2021-02-24T06:53:34-0500

The possible number of tails is: 0,1,2,3,4. Those are the values that "Y" can take.

In order to compute the respective probabilities we will use the binomial distribution. We assume that the probability of a tail and the probability of a head are equal. We receive:

"P(Y=0)=\left(\frac{1}{2}\right)^4=\frac{1}{16};P(Y=1)=C_4^1\left(\frac{1}{2}\right)^4=4\left(\frac{1}{2}\right)^4=\frac{1}{4};"

"P(Y=2)=C_4^2\left(\frac{1}{2}\right)^4=\frac{3\cdot4}{2}\left(\frac{1}{2}\right)^4=\frac{3}{8}; P(Y=3)=C_4^3\left(\frac{1}{2}\right)^4=\frac{1}{4}";

"P(Y=4)=C_4^4\left(\frac{1}{2}\right)^4=\frac{1}{16}"

We check that the sum of all probabilities is 1: "P(Y=0)+P(Y=1)+P(Y=2)+P(Y=3)+P(Y=4)="

"\frac{1}{16}+\frac{1}{16}+\frac{1}{4}+\frac{1}{4}+\frac{3}{8}=\frac{1}{2}+\frac{1}{2}=1"